OpenStudy (anonymous):

Anyone here good with differentiating functions based on the derivative rules?

OpenStudy (amistre64):

if only there werent so many derivative rules ....

OpenStudy (anonymous):

i bet you have everyone's rapt attention now

OpenStudy (amistre64):

survey says? YES! Ding!!

OpenStudy (anonymous):

i think he has been waiting for a "yes"

OpenStudy (anonymous):

@amistre, i going to turning early tonight, later

OpenStudy (anonymous):

f(x)= square root of 30 would equal 0 because of d/dx[x]=0 rule right?

OpenStudy (amistre64):

Ciao :)

OpenStudy (anonymous):

by x i meant c

OpenStudy (amistre64):

yes

OpenStudy (amistre64):

sqrt(30) is a constant

OpenStudy (amistre64):

the derivative of a constant is ....0 since it never changes

OpenStudy (amistre64):

[C]' = 0 think of it as the power rule of x^0

OpenStudy (anonymous):

they are trying to get you to say it is \[\frac{1}{2\sqrt{30}}\] so then they can laugh at how they tricked you

OpenStudy (anonymous):

so it isnt 0?

OpenStudy (amistre64):

is your question; what is the derivative of sqrt(30) ?

OpenStudy (anonymous):

yes it is zero.

OpenStudy (amistre64):

f(x)= sqrt(30) would equal 0 because of d/dx[C]=0

OpenStudy (anonymous):

what about differentiate the function f(x)=x^3-4x+6 d/dx[x^n]=nx^(n-1) apply right?

OpenStudy (amistre64):

yes, its just a poly that derives out by the power rule

OpenStudy (amistre64):

and d/dx[6]=0

OpenStudy (anonymous):

can you be elborate, im kind of confused